### v1

# egoPPG: Heart Rate Estimation from Eye-Tracking Cameras in Egocentric Systems to Benefit Downstream Vision Tasks

Abstract
Egocentric vision systems aim to understand the spatial surroundings and the wearer’s behavior inside it, including motions, activities, and interaction with objects. Since a person’s attention and situational responses are influenced by their physiological state, egocentric systems must also detect this state for better context awareness. In this paper, we propose egoPPG, a novel task for egocentric vision systems to extract a person’s heart rate (HR) as a key indicator of the wearer’s physiological state from the system’s built-in sensors (e.g., eye tracking videos). We then propose EgoPulseFormer, a method that solely takes eye-tracking video as input to estimate a person’s photoplethysmogram (PPG) from areas around the eyes to track HR values—without requiring additional or dedicated hardware. We demonstrate the downstream benefit of EgoPulseFormer on EgoExo4D, where we find that augmenting existing models with tracked HR values improves proficiency estimation by 14%. To train and validate EgoPulseFormer, we collected a dataset of 13+ hours of eye-tracking videos from Project Aria and contact-based blood volume pulse signals as well as an electrocardiogram (ECG) for ground-truth HR values. 25 participants performed diverse everyday activities such as office work, cooking, dancing, and exercising, which induced significant natural motion and HR variation (44–164 bpm). Our model robustly estimates HR (MAE=8.82 bpm) and captures patterns (r=0.81). Our results show how egocentric systems may unify environmental and physiological tracking to better understand user actions and internal states.
Figure 1: Implementing a novel vision task egoPPG, EgoPulseFormer tracks a person’s heart rate (HR) from eye-tracking videos captured by unmodified egocentric vision headsets. Our method estimates the person’s photoplethysmogram (PPG) of the blood volume pulse from areas around the eyes to extract HR values. For training and validation, we collected egoPPG-DB, a dataset of participants’ eye-tracking videos during everyday activities with synchronized ground-truth PPG signals (white contact sensor) and HR values (ECG chest strap).
1 Introduction
Egocentric vision systems, such as Mixed Reality (MR) glasses by Meta [56], Magic Leap [46], and others have emerged as powerful devices for capturing and analyzing a person’s behavior as well as their surrounding environment from a first-person perspective. The wider availability of promising wearable capture platforms (e.g., Project Aria glasses [23]) has sparked a large amount of research on egocentric vision tasks for environment understanding and navigation, including localization [71, 41, 74], and simultaneous localization and mapping (SLAM) [17, 38, 68]. Since egocentric systems can simultaneously capture parts of the wearer’s behavior, many efforts have investigated egocentric action recognition [52, 89, 94, 44, 86] and hand-object interaction [73, 32, 95, 22] to understand user behavior. Large-scale datasets have also been introduced in recent years to accelerate data-driven research in this domain, such as Ego4D [32], Nymeria [51], and EgoExo4D [33], offering rich, multimodal data to enable training and evaluation for these tasks.
In addition to spatial awareness, understanding the user’s behavior, their attention, and intent is equally important for egocentric systems [88, 3, 58]. For instance, anticipating the user’s next action is crucial for applications in navigation, personalized feedback, and autonomous assistance [20, 91, 84]. Objects of interest are commonly estimated from analyzed gaze patterns [27, 47, 37, 45] to support behavioral analysis and social understanding [26, 40, 32, 33].
However, holistically modeling a person’s behavior and intent requires knowledge of their physiological state, which influences cognitive performance, attention, and situational responses [25, 15, 53, 81, 76]. Key components of physiological state include cardiovascular indicators such as heart rate (HR) and electrodermal activity [59, 79, 1, 9], which reflect emotions, stress, fatigue, and alertness [65, 11, 59, 79, 1]. Capturing these dynamics can thus benefit models of human behavior to enable a richer understanding of user actions for adaptive systems.
In this paper, we introduce egoPPG, a novel task for egocentric vision systems to accurately extract a person’s HR measurements from the system’s built-in sensors, specifically the eye-tracking cameras in unmodified headsets. We then propose EgoPulseFormer, a novel method that implements egoPPG on Project Aria glasses to demonstrate the benefit of egoPPG for existing downstream vision tasks on large egocentric datasets. Our learning-based method EgoPulseFormer is designed to recover the person’s photoplethysmogram (PPG) from the subtle fluctuations in skin intensity due to light absorption in pulsatile arteries beneath the skin surface following a blood volume pulse (BVP), in particular deriving it from regions around the wearer’s eye for robust tracking. Leveraging the infrared (IR) illuminant of the eye tracker, it work across light conditions [50]. In addition, we demonstrate the benefits of estimated HR values in egocentric recordings for a key vision task downstream: We augment an existing architecture with EgoPulseFormer and show its impact on EgoExo4D’s proficiency estimation benchmark, whose accuracy improves by 14.1% with EgoPulseFormer’s HR estimates.
The sensing configuration of our task can be considered a hybrid between typical contact-based BVP sensors (e.g., those in smartwatches [5, 31, 63, 21, 13, 61]) and rPPG methods that aim to extract HR from a person’s face using a camera [39, 82, 66]. While egocentric glasses are body-worn much like wrist watches, they couple more loosely to the body and are subject to considerable motion artifacts. Unlike contact sensors, eye trackers observe the wearer’s eye regions from a short distance and capture eye motions and blinks, leading to ambiguity and noise for capturing fluctuations in skin intensity. Unlike rPPG configurations, egocentric capture systems move with the wearer’s body and head, and eye trackers use controlled illumination.
Therefore, we designed EgoPulseFormer to extract BVP from the region with the least motion—around the wearer’s eyes—by incorporating spatial attention within our model’s backbone. We validate EgoPulseFormer’s efficacy on a novel dataset that we collected to capture some of the activities included in large-scale egocentric datasets alongside physiological reference recordings. Our dataset egoPPG-DB contains 13 hours of recordings from 25 participants, who wore Project Aria glasses and performed five real-world tasks with varying motion and intensity, causing their HR values to reach levels between 44–164 bpm.
We summarize our key contributions as follows:
  1. 1.
    egoPPG as a novel task and EgoPulseFormer as an HR estimation method for egocentric systems that operates on eye-tracking videos. Our method robustly predicts continuous HR across a series of activities and interactions (MAE=8.82 bpm), with a 27% lower error than current state-of-the-art rPPG models [12, 92, 48, 90].
  2. 2.
    egoPPG-DB, a dataset of eye-tracking videos and synchronized BVP (contact-based) and ECG recordings (chest strap-based) to verify all physiological signals. We captured these across diverse everyday activities that were inspired by those included in existing large-scale egocentric datasets [32, 51, 33].
  3. 3.
    a validation of egoPPG’s downstream benefits for proficiency estimation. Augmenting EgoExo4D with continuous HR values and additionally feeding them into an existing architecture, we demonstrate the implications of our method on the proficiency estimation benchmark with an increase in accuracy by 14.1%.
2 Related work
Egocentric vision. In recent years, research in egocentric vision has surged, driven by advances in AR/VR glasses [4, 36, 46, 56, 57, 23], which provide new ways for understanding user interaction from a first-person perspective. Much of this work has focused on tasks such as action recognition [52, 89, 94, 44, 86] and anticipation [30, 16, 86, 60], full-body pose estimation [75, 87], responding to user needs [69, 67, 91], and social behavior analysis [26, 40, 32]. Additionally, tracking vital signs in AR/VR settings and for affective computing applications [65, 11, 59, 79, 1] has become an important tool for understanding users’ physiological states [54], further aiding in understanding users’ behavior, their attention, and intent [88, 3, 58].
Physiological measurements. Wearable sensors have had a tremendous impact on health monitoring in recent years, enabling continuous measurement of key physiological metrics, such as heart rate (HR), oxygen saturation, and activity levels [63, 21, 13, 62]. Heart rate (HR), in particular, is a key measure for assessing an individual’s health and performance [29, 24, 72, 42]. While wearable sensors, such as wrist-worn smartwatches, provide accurate HR measurements and also challenging scenarios (e.g., exercising), they are intrusive and can cause discomfort [43]. Recent research has, thus, extensively explored using cameras as an unobtrusive, non-contact alternative for measuring HR, generally called remote photoplethysmography (rPPG) [39, 82, 66]. rPPG measures HR based on subtle color changes in the skin caused by the BVP. Generally, rPPG methods can be broadly divided into traditional signal processing techniques [39, 82, 66, 8, 18, 19, 83] and deep learning-based approaches [12, 92, 48, 10, 90]. So far, rPPG has been mostly applied to facial videos with the camera and user being stationary, such as while sitting in front of a laptop, as it requires a continuous video feed of the same skin region. This limitation is shown in current rPPG datasets, which primarily capture individuals in seated positions with either a stationary camera directed at their face [78, 34, 7, 64, 70] or requiring users to hold a smartphone steadily in front of their face [80]. As a result, rPPG is not feasible to be deployed in more dynamic settings, such as during exercise.
Eye tracking cameras. Eye tracking in egocentric vision systems is mostly done using inward-facing cameras direct at the eyes [2]. Even during motion, eye tracking in VR devices demonstrated accurate performance showcasing that the cameras remain almost stationary relative to the user’s eyes [14]. Furthermore, IR illumination makes them robust to lighting variations and low-light conditions [50]. To the best of our knowledge, videos from eye tracking cameras have not yet been explored for HR estimation using the BVP despite their promise to enable unobtrusive HR measurements during everyday life.
3 Overview
Our aim is to enable egocentric vision systems i) to track a person’s physiological state via continuously estimated HR and ii) to integrate these HR estimates into downstream tasks that benefit from knowledge of the user’s state. Sec. 4 first describes our dataset of synchronized eye-tracking videos and ground-truth HR measurements during a series of everyday activities. Sec. 5 then outlines our method EgoPulseFormer that continuously estimates a person’s BVP from eye-tracking videos and derives HR values from it. Fig. 3 illustrates our approach. To demonstrate EgoPulseFormer’s benefit for downstream applications, we leverage HR estimation for the user proficiency benchmark of the EgoExo4D dataset, described in Sec. 6. Finally, Sec. 7 provides all results from our evaluations and Sec. 8 discusses our findings.
4 egoPPG-DB
The egoPPG-DB dataset was developed to support HR estimation from eye-tracking videos under real-world conditions, with a protocol designed to elicit significant motion and fluctuations in HR. By including diverse everyday activities, we provide a challenging benchmark for egocentric HR estimation models.
4.1 Recruiting and recording
We recruited $N=25$ participants (12 female, 13 male, ages 19–32, $\mu=25.1$ and $\sigma=3.3$) on a voluntary basis, resulting in over 13  hours of video recordings. Based on the Fitzpatrick scale [28], 9 participants had skin type II, 10 had skin type III, 2 had skin type IV, and 4 had skin type V. All participants signed a consent form before the data collection, agreeing with using and sharing their data for academic and non-commercial purposes. Participants were instructed to avoid wearing makeup prior to the recording. The data collection was approved by the anonymized Ethics Commission (no. #anonymized). In terms of dataset size by duration, egoPPG-DB is third amongst the longest rPPG datasets [78, 34, 7, 64, 70, 80] as listed in Tab. 6.
4.2 Apparatus
Fig. 2 illustrates our experimental setup. We used Project Aria glasses [23] with Profile 21 to record eye tracking videos at 30 fps with a resolution of $320\times 240$ pixels per eye. To capture ground truth PPG measurements, with which we train our model, we developed a custom sensor that records PPG data offline at 128 Hz. The sensor consists of a main board, mounted on the left side of the frame, featuring a DA14695 system-on-chip interfacing with a MAX86141ENP+ PPG sensor. The LEDs and photodiodes used by the PPG sensor are embedded in the left nose pad and connected to the main board using a flat flexible cable. For each participant, we individually adjusted the nose pad position to ensure the sensor aligned with their left angular artery [35]. To validate our custom PPG sensor, we also recorded gold-standard ECG data using a movisens ECGMove 4 chest belt sampling at 1024 Hz. We synchronized all devices at the start and end of each recording with a synchronization pattern, using their built-in IMUs.
Figure 2: Apparatus used to record the egoPPG-DB dataset.
4.3 Capture protocol
The average duration of the recording of participant was 32 minutes. The capture protocol comprised 5 activities (Tab. 1): watching a video, office work, kitchen work, dancing, and exercising on an indoor bike (Fig. 1). We included these activities for three purposes: (1) Incorporate everyday activities including the corresponding HR changes and motion artifacts; (2) cover a wide range of HR values (low HR when watching a video vs. high HR when exercising), and (3) resemble activities that were captured in large-scale egocentric vision datasets, such as EgoExo4D [33] or Nymeria [51]. In Tab. 8, we give a detailed description of each activity, and in Tab. 3 we show the mean HR values for each activity. Exercising on the bike produced the highest HR values (113 bpm), whereas watching the video resulted in the lowest mean HR (71 bpm).
4.4 Dataset and signal quality verification
To evaluate that the contact PPG sensor, whose signal we later use as the target for model training, produces accurate HR values, we evaluate it against the gold-standard ECG. We assessed the performance by calculating the MAE and Pearson correlation between HR estimates from the ECG and PPG signals for each participant using a 30-second sliding window. For activity labeling, we manually annotated the start and end times of each task (see Tab. 1) for each participant using the Point of View (POV) RGB videos recorded by the Project Aria glasses. To ensure that the signal quality of the contact PPG is sufficient for model training, we excluded all tasks with an MAE over 3.0 bpm between the PPG and ECG, which can happen when the PPG sensor occasionally loses alignment with the angular artery due to movement. This applied to 20 of 150 tasks (13%, see Tab. 7). During the remaining tasks, our custom-built PPG nose sensor achieved very high accuracy, with an average MAE of 1.3 bpm and a mean correlation of 0.94 compared to the ECG signal, showing its suitability as ground truth.
activity actions minutes
Watch video Watch a documentary 5
Office work Work on a computer 4
Write on a paper 2
Talk to the experimenter 2
Walking Walk to the kitchen 1
Kitchen work Cut vegetables 5
Prepare a sandwich
Wash the dishes
Walking Walk to the dancing room 1.5
Dancing Follow random dance video 5
Exercise bike Ride an exercise bike 5
Walking Walk back to the start 1.5
Table 1: Capture protocol for recording the egoPPG-DB dataset.
5 EgoPulseFormer: a first method for egoPPG
5.1 Problem definition
Our objective is to estimate BVP and HR from periodic changes in pixel intensity in eye-tracking video frames $\boldsymbol{F}\in\mathbb{R}^{w\times h}$. Physically, this means extracting a physiological signals from the information in the light reflected by the arteries and arterioles that carry blood beneath the skin. This light reflection can be modeled as a combination of diffuse and specular reflections. Wang et al. [83] model the reflected light intensity $C(t)$ as:
$\boldsymbol{C(t)}=I(t)(\boldsymbol{v_{s}}(t)+\boldsymbol{v_{d}}(t))+\boldsymbol{v_{n}}(t)$ (1)
where $I(t)$ is the luminance intensity, $\boldsymbol{v_{s}}(t)$ the specular reflection, $\boldsymbol{v_{d}}(t)$ the diffuse reflection, and $\boldsymbol{v_{n}(t)}$ the sensor noise. While the specular reflection $\boldsymbol{v_{s}}(t)$ lacks pulsatile information, the diffuse reflection $\boldsymbol{v_{d}}(t)$ contains information about the absorption and scattering of the light in skin tissue[83]. Thus, $\boldsymbol{v_{d}}(t)$ can be further decomposed as:
$\boldsymbol{v_{d}}(t)=\boldsymbol{u_{d}}d_{0}+\boldsymbol{u_{p}}p(t)$ (2)
where $\boldsymbol{u_{d}}$ is the unit color vector of the skin, $d_{0}$ the stationary reflection strength, $\boldsymbol{u_{p}}$ the relative absorption, and $p(t)$ the signals of interest. $p(t)$ is in our case the BVP, which our model aims to learn from the camera recordings.
Figure 3: Architecture of our model for continuous BVP estimation from eye-tracking videos and consecutive HR computation.
5.2 Deep learning model
Our architecture is built upon a 3D CNN backbone [92] with a temporal input length of $T=128$ frames (corresponding to 4.3 seconds) downsampled to $(h=48)\times(w=128)$ pixels, resulting in an input of dimensions $(T,C,w,h)$. The channel is $C=1$ in our case, as our input is from monochrome videos. Compared to the video of a person’s face, which is usually used for rPPG tasks, eye tracking videos offer additional challenges. While the bulbar conjunctiva (white of the eyes) contains many blood vessels from which the BVP could theoretically be estimated, eyes typically move strongly during everyday situations and are closed while blinking (see participant 2 in Fig. 4). Consequently, extracting the BVP from the eye regions would introduce substantial motion artifacts and reduce the signal-to-noise ratio (SNR). In contrast, when qualitatively analyzing eye tracking images, we see that the skin around the eyes exhibits considerably less motion than the eyes themselves and could thus provide a more stable source of BVP information. To address this, we introduce spatial attention modules [85] before each pooling (see Fig. 3) to allow our network to focus on high-SNR regions, such as the skin, and reduce the influence of low-SNR regions with frequent motion, like the eyes. Given an input feature map $\boldsymbol{F}\in\mathbb{R}^{T\times C\times w\times h}$, the spatial attention modules infers a spatial attention map $\boldsymbol{M_{s}}\in\mathbb{R}^{T\times 1\times w\times h}$ as:
$\boldsymbol{M_{s}}(\boldsymbol{F})=\sigma*(f^{7\times 7}([\boldsymbol{F_{avg}};\boldsymbol{F_{max}}]))$ (3)
where $\sigma$ is the sigmoid function, $f^{7\times 7}$ a $7\times 7$ convolution operation and $\boldsymbol{F_{avg}}\in\mathbb{R}^{T\times 1\times w\times h}$ and $\boldsymbol{F_{max}}\in\mathbb{R}^{T\times 1\times w\times h}$ are the average-pooled and max-pooled feature maps respectively. The final output $\mathbf{F_{out}}$ of the attention process is then:
$\mathbf{F_{out}}=\mathbf{M_{s}}\otimes\mathbf{F},$ (4)
Furthermore, individual variations in the fit of the glasses result in different parts of the skin around the eyes being visible. For some individuals, the eye tracking cameras capture only the areas above the eyes, for others, only below, and in some cases, the glasses sit at an incline (see Fig. 4). To account for such variations, we apply three targeted data augmentations during training that reflect these specific differences in camera angles and coverage: (1) random rotation between -20 and +20 degrees to account for slight inclinations in the glasses’ positioning; (2) random horizontal cropping to help the network distinguish between high and low SNR regions across various skin areas and camera positions; and (3) horizontal and vertical flipping to further increase robustness to individual differences in skin region visibility. Finally, to help the network focus on the changes between the frames caused by the periodic BVP, we use the standardized consecutive frame differences of the eye tracking videos as input into our network. We standardize each frame by subtracting the mean pixel intensity and dividing it by the standard deviation of the pixel intensities values [12]. We use the standardized consecutive differences of the PPG signals from the nose as the labels for our model. Fig. 3 shows our architecture with the input preprocessing demonstrated on the left side, and an example learned spatial attention map on the right side. For one batch, the total number of FLOPS for our model is about 164 GigaFLOPS. The total number of parameters is about 770k, where the spatial attention modules added only 400 parameters.
Figure 4: Left: Head geometry determines the regions that the eye tracker captures. Right: Learned spatial attention maps show that eye regions are excluded and EgoPulseFormer instead extracts BVP from the surrounding skin regions, which moves less than the eyes.
5.3 Experiments setup
5.3.1 Training
We trained all implemented models using five-fold cross-validation split by participants to ensure a strict separation between training, validation, and test sets. We iteratively held out the data from five participants (20%) as the test set, two as validation, and used the remaining participants as the training set. The training was conducted with a batch size of 4 for 100 epochs, a learning rate of 0.0009, and mean squared error (MSE) as the loss function. In addition to our model, we used four baseline networks (one state-space and three CNNs), which have state-of-the-art performance for rPPG, to compare the performance of our proposed model to the performance of these established models. Our model was trained on a GeForce RTX 4090, with a total runtime of about 15 hours for all folds.
5.3.2 Metrics
To assess model accuracy, we use the mean absolute error (MAE), root mean squared error (RMSE), mean absolute percentage error (MAPE), and Pearson correlation (r) using a non-overlapping 60-second sliding window, as commonly used for rPPG [18, 92, 48, 55, 49].
5.3.3 Video sampling rate
While we recorded the eye tracking videos with 30 fps, large-scale datasets such as EgoExo4D [33] or Nymeria [51] used only 10 fps. To assess the impact of reduced frame rates, we evaluated i) model performance when downsampling our videos to 10 fps by retaining only every third frame and ii) model performance when downsampling to 10 fps, then linearly interpolating between frames to upsample to 30 fps again. For both scenarios, we train from random initialization.
5.3.4 Hyperparameter optimization
We also provide insights into optimal hyperparameter configurations by analyzing the impact of image size, input length, loss function, and choice of label (see Tab. 9).
6 Downstream use for proficiency estimation
To demonstrate the utility of predicting a user’s physiological state for egocentric vision applications, we use the user proficiency estimation benchmark from the EgoExo4D dataset, which contains over 5000 videos from 740 participants performing skilled human activities [33]. The user proficiency estimation benchmark aims to classify the proficiency of a user (novice, early expert, intermediate expert, late expert) using only egocentric video clips (Ego), only exocentric video clips (Exo), or all video clips together (Ego + Exo). Our goal was to assess if we can improve the performance of the current baseline model (TimeSFormer [6]) when integrating our predicted HR data into the network. This results in two additional configurations: using egocentric videos and HR together (Ego + HR), and using all video clips and HR together (Ego + Exo + HR). To predict the continuous HRs for all EgoExo4D videos, we use our proposed method, pre-trained on egoPPG-DB.
We implement the TimeSFormer model in exactly the same configuration as for the current benchmark results [33] with a clip size of 16 frames and a sampling rate of 16, trained for 15 epochs on four GeForce RTX 4090. We use all videos of the EgoExo4D dataset, for which the proficiency estimation labels are available (using the official benchmark training and validation sets) and which have at least 16 frames at a sampling rate of 16, resulting in totally 2044 videos. From the official training set, we use 10% as validation, and the held-out official validation set for testing. We summarize our predicted HR data by calculating five features (mean, standard deviation, minimum and maximum HR, and mean HR change) for the corresponding videos. To integrate our HR predictions into the TimeSFormer architecture, we simply feed our calculated features into a 50-parameter linear layer and concatenate the output with the backbone’s output before feeding it into the network head. We train all models from random initialization and evaluate proficiency estimation using top-1 accuracy per the EgoExo4D protocol.
7 Experiments
7.1 Heart rate estimation
7.1.1 Baseline
We employed signal processing to verify that the BVP signal is present in the eye tracking videos, to determine in which regions the SNR is highest, and to establish a baseline for comparison. Since the glasses remain mostly stable throughout the recording, we manually define two spatial cropping regions per participant. One region that includes mostly skin, and one region that includes mainly eyes (see Fig. 4). After spatially cropping the images, we calculate the mean pixel intensity values and remove motion artifacts by discarding any changes in the signal that are outside the first and third quantiles of the signal’s changes. Finally, we filter the signal with a 4th order Butterworth bandpass filter between 0.6 and 3.0 Hz (corresponding to 36 to 180 bpm) to obtain the BVP (see Fig. 6).
7.1.2 EgoPulseFormer: an egoPPG method
Using our proposed network, we obtain an MAE of 8.82 bpm and a correlation of 0.81 between our predicted HR and the ground truth HR (see Tab. 2). This is an improvement of 3.27 bpm (27.0%) of the MAE and 0.15 for the correlation compared to the second-best network (PhysNet [92]). Split by activity, we obtain the lowest MAE while the participants are watching a video (MAE of 6.25 bpm) and the highest MAE during exercising on a bike (MAE of 13.51 bpm) and while dancing (MAE of 10.02 bpm), which are both also the tasks with the highest normalized motion magnitude. We define the motion magnitude as the root mean squared sum of the absolute differences across the 3-axis IMU recorded by the Aria glasses. We, then, normalize it between zero and one across all activities to get a measure of the amount of motion of each activity. Furthermore, using simple signal processing and spatial cropping, we obtain an MAE of 12.40 and a correlation of 0.50 when using the skin region around the eyes. When using the eyes regions as input, the MAE increases to 14.60, and the correlation drops to 0.20. This is also reflected in spatial attention maps that our model implicitly learns (see Fig. 4), which exclude the eyes for predicting the HR. To qualitatively cross-check these results, Fig. 6 shows an example plot of the raw mean intensity values (before filtering) of the skin region compared to the eye region with the BVP clearly visible for the skin region. Tab. 4 shows the results when downsampling our videos to 10 fps. The MAE increases to 12.40 bpm and the correlation decreases to 0.52 when training and testing using 10 fps. When upsampling the videos again to 30 fps using linear interpolation, the MAE decreases to 10.20 bpm and the correlation increases to 0.77.
Model MAE RMSE MAPE r
DeepPhys [12] 28.26 31.97 36.68 0.08
TS-CAN [48] 26.32 32.39 29.13 0.11
Mean intensity eyes 14.60 18.18 18.37 0.20
PhysMamba [90] 13.94 16.86 17.76 0.61
Mean intensity skin 12.40 15.54 15.29 0.50
PhysNet [92] 12.09 15.43 15.14 0.66
EgoPulseFormer (ours) 8.82 12.03 10.82 0.81
Improvement over -3.27 -3.4 -4.32 +0.15
second-best method
Table 2: Results for HR prediction from eye tracking videos using different models (EgoPulseFormer and established rPPG baselines).
Activity $\mu$ HR Motion MAE RMSE MAPE
magnitude
Video 71.45 0 6.25 8.50 9.88
Office 75.65 0.45 8.90 11.97 12.60
Kitchen 85.31 0.54 8.75 11.11 10.42
Dancing 89.07 1.00 10.02 12.79 11.22
Bike 113.06 0.77 13.51 16.52 11.04
Walking 93.71 0.30 7.02 9.31 6.70
Table 3: Results for HR prediction split by activity.
Input video MAE RMSE MAPE r
10 fps (other datasets) 12.40 16.80 14.10 0.52
Upsampled to 30 fps 10.20 13.56 12.78 0.77
Table 4: Results for HR prediction with different frame rates. In the first row, we downsample our videos to a frame rate of 10 fps, commonly used by large-scale datasets such as EgoExo4D [33]. In the second row, we first downsample our videos to 10 fps and then upsample them to 30 fps by linearly interpolating between frames.
7.2 Downstream task: proficiency estimation
Tab. 5 summarizes the results of our experiments to evaluate the value of HR estimation for the proficiency estimation benchmark on the EgoExo4D dataset. We see that integrating our predicted HRs into the TimeSFormer model [6] improved accuracy for all scenarios but one. Additionally, we also achieved the highest accuracy for each individual scenario with our HR integration for the same scenarios. When combining the egocentric videos with our predicted HRs, we achieved an overall accuracy of 45.20%, a 14.1% increase compared to using egocentric videos alone. The largest gains appeared in the cooking and dancing tasks, where accuracy rose from 20.00% to 40.00% and from 43.44% to 53.27%, respectively. Also, when using the egocentric videos, exocentric videos, and our predicted HRs together, the accuracy increased by 8.64% from 39.00% to 42.37% compared to using only the egocentric and exocentric videos. Using only exocentric videos yielded the lowest overall accuracy at 35.93%.
Scenario Majority Ego Ego + HR (ours) Exo Exo + HR (ours) Ego + Exo Ego + Exo + HR (ours)
Basketball 38.00 45.45 47.47 49.24 49.24 49.49 52.52
Cooking 0.00 20.00 40.00 33.75 40.00 25.00 40.00
Dancing 24.59 43.44 53.27 45.08 48.36 50.82 55.73
Music 57.89 78.94 81.58 57.89 57.89 57.89 60.53
Bouldering 15.29 24.50 27.81 10.26 12.58 15.89 18.54
Soccer 62.50 50.00 56.25 76.56 75.00 75.0 62.50
Overall 27.80 39.69 45.29 35.93 35.93 39.00 42.37
Table 5: Results for proficiency estimation benchmark on EgoExo4D dataset. Note that for all scenarios except Soccer, the accuracy increases when integrating EgoPulseFormer’s heart rate estimate into the existing and otherwise unmodified baseline model.
8 Discussion
8.1 Heart rate estimation
When evaluating our proposed method on egoPPG-DB, we showed that HR can reliably be predicted from eye tracking videos of unmodified egocentric vision headsets. Compared to the performances achieved on popular rPPG datasets such as PURE [78], UBFC-RPPG [7], and MMPD [80], we achieved very competitive performance in challenging settings [49]. While the lowest reported MAE for UBFC-RPPG in the rPPG-toolbox [49] is 1.21 bpm, its participants sit almost motion-free with their eyes closed. Already on MMPD, a dataset with varied lighting and little motion, the lowest reported MAE increases to 10.23 bpm [49], as it incorporates more realistic challenges due to recordings on mobile devices with light motion (head rotation, talking, and taking selfies). In comparison, using EgoPulseFormer, we achieved a lower MAE of 8.82 bpm while even having tasks with heavy motion (dancing, walking stairs) and very high HR changes (44–164 bpm).
8.1.1 Performance depending on activity
Analyzing our results split by activity (see Tab. 3), we see that the tasks with the highest mean HR (bike, 113 bpm) and motion magnitude (dancing) also have the highest MAE. Given the substantially bigger motion artifacts and HR variability (HRs between 44 and 164 bpm) in our dataset — captured during diverse everyday activities and physical exercises — we believe that our results demonstrate the robustness of EgoPulseFormer in more dynamic, everyday conditions.
8.1.2 Performance depending on method
For our model, we leveraged spatial attention maps to improve performance. When qualitatively analyzing the learned spatial attention maps, we see that our model implicitly learned to exclude the eyes for estimating the BVP from the eye tracking videos (see Fig. 4). This aligns with our obtained qualitative (see Fig. 6) and quantitative (see Tab. 2) results using simple signal processing, which show a higher SNR for the skin region compared to the eyes. When analyzing the performance of other state-of-the-art rPPG models on our dataset, we found that only PhysNet [92] and PhysMamba [90] achieved reasonable performances, though their results did not reach our performance.
8.1.3 Performance depending on camera fps
Using eye tracking videos recorded at only 10 fps considerably decreases performance (see Tab. 4). However, upsampling the frame rate to 30 fps through linear interpolation between frames substantially improves the performance again. This is especially important as many large-scale datasets, such as EgoExo4D [33] or Nymeria [51], for which predicting a user’s physiological state could help for further downstream tasks, are recorded at only 10 fps.
8.2 Benefits for proficiency estimation downstream
We found that incorporating HR data into the baseline model of the proficiency estimation task substantially improved accuracy across both configurations. The egocentric videos combined with the HR achieved the highest overall accuracy at 45.29%, marking a 14.1% increase over using only egocentric videos (39.69%). Adding HR especially improved accuracy for cooking (from 20% to 40%) and dancing (from 43.4% to 53.3%), which had the lowest accuracies besides bouldering when using only egocentric videos, demonstrating the value of HR in enhancing model performance. Combining egocentric videos, exocentric videos, and HR provided further accuracy gains for some scenarios, achieving the best results for basketball, cooking, and dancing. Results using exocentric views alone were lower overall, which is consistent with benchmark results [33]. These findings suggest that adding HR data as an auxiliary signal enhances performance for the proficiency estimation benchmark on the EgoExo4D dataset. For training and testing, we used the available subset of EgoExo4D videos for which proficiency labels are available, following the official training and validation splits. While our used data shows slight variations from the official release in majority class distributions and accuracy scores, the observed trends align well with the established benchmark results.
8.3 Limitations and future work
We observed the highest MAE in tasks with elevated HRs and motion, such as dancing and biking. We believe that promising approaches to address these limitations could be to record more tasks with high HRs and integrating the IMU data from the glasses. Furthermore, future studies could extend data collection to outdoor settings to assess the impact of varying lighting conditions. While we ensured a balanced gender ratio (13 male, 12 female), our sample size of 25 participants restricts broader demographic conclusions. In future work, we aim to expand the dataset to investigate performance variations across different age groups, skin types, and ethnicities. Finally, we believe that it is a highly interesting problem to explore further downstream applications of a user’s physiological state, such as personalized feedback, autonomous assistance, as well as health-related applications.
9 Conclusion
We introduced egoPPG, a novel task for egocentric vision systems to extract the wearer’s heart rate for integrating their physiological state into egocentric vision tasks downstream. Our method EgoPulseFormer processes input from the eye tracking cameras on unmodified egocentric vision systems to robustly estimate the person’s HR in various everyday scenarios. We validate EgoPulseFormer’s robustness on our dataset egoPPG-DB and demonstrate significant improvements over existing rPPG models. With HR estimations from EgoPulseFormer we significantly improve the proficiency estimation benchmark on the large-scale EgoExo4D dataset. Our results emphasize the potential of physiological insights obtained via egoPPG methods for further egocentric vision applications.
Supplementary Material
10 Related datasets
Tab. 6 gives a comprehensive comparison of the dataset size and activities of related remote photoplethysmography (rPPG) datasets. In terms of hours of recordings and recorded frames, egoPPG-DB is the third largest dataset. Furthermore, we see that all comparable rPPG datasets only include activities with very little motion and heart rate (HR) changes such as watching videos, head rotations or talking. In contrast, egoPPG-DB features a wide variety of challenging everyday activities, such as kitchen work, dancing riding and exercise bike, which induce significant motion artifacts and HR changes.
Dataset Part. Frames Hours Tasks
PURE [78] 10 110 K 1 Resting, talking, small head movements
MAHNOB-HCI [77] 27 2.6 M 12 Watching videos
MMPD [80] 33 1.2 M 11 Resting, head rotation, selfie videos
MMSE-HR [93] 40 310 K 2 Talking, watching videos, experiencing different emotions
UBFC-rPPG [7] 43 150 K 1.5 Gaming on a computer
UBFC-PHYS [70] 56 2.4 M 19 Resting, Trier Social Stress Test
VIPL-HR [64] 107 4.3 M 20 Resting, talking, head rotation, different lighting conditions
egoPPG-DB (ours) 25 1.4 M 13 Watching videos, office and kitchen work, dancing, biking, walking
Table 6: Summary of existing datasets for rPPG.
11 Excluded tasks
For all participants and activities, we checked the mean absolute error (MAE) between the predicted HR from our custom contact PPG sensor on the nose and the gold standard ECG from the chest belt. We excluded all tasks with an MAE over 3.0 beats per minute (bpm), which can happen, for example, when the PPG sensor loses alignment with the angular artery due to movement. In this way, we ensured that the photoplethysmography (PPG) signal signal from the nose, which we used as the target signal to train our model, is highly accurate. As a result, we had to exclude 20 out of the 150 tasks (13%), which we list in Tab. 7. We can see that this applied only to tasks with more motion (dancing, exercise bike, and walking). Since the participants had to walk multiple stairs throughout the data recording, this mostly happened during walking.
Activity Excluded participants
Watch video
Office work
Kitchen work
Dancing 012, 015
Exercise bike 009, 012, 014, 015, 016, 023
Walking 004, 012, 013, 014, 018, 021, 022
Table 7: Detailed table of all excluded tasks.
12 Detailed description of activities
Tab. 8 gives a comprehensive description of the actions for each activity during our recording. Generally, participants were free to talk during the entire duration of the recording and conduct the tasks as they would do it normally. For example, during the kitchen work, the participants were completely free to prepare the sandwich and if they would like to eat or drink while doing it.
Activity Actions Description
Watch video Watch a documentary Watch a relaxing documentary on a computer.
Office work Work on a computer Randomly browse through websites and type text from a PDF into Word.
Write on a paper Write a text from a PDF on a computer onto a piece of paper.
Talk to the experimenter Have a free, unscripted conversation with the experimenter.
Walking Walk to the kitchen Walk along a hallway, down the stairs into the kitchen.
Kitchen work Get ingredients Get all ingredients for a sandwich from the fridge.
Cut vegetables Get a cutting board, knife and a plate and cut vegetables.
Prepare a sandwich Put the bread into the toaster and afterward freely prepare sandwich.
Eat sandwich/drink Participants are free to eat the sandwich or drink during the recording.
Wash the dishes Wash everything used while preparing the sandwich.
Walking Walk to the dancing room Walking along a hallway into a new room for dancing and biking.
Dancing Follow random dance video Choose a dance video and afterward follow it.
Exercise bike Ride an exercise bike Ride an exercise bike with moderate to high intensity.
Walking Walk back to the start Walk back to the start either up the stairs or using the elevator.
Table 8: Detailed capture protocol and action descriptions of the egoPPG-DB dataset.
13 Data recording
In Fig. 5, we show a variety of different images and people of our data recording from a third person view to visualize the apparatus and capture protocol. All participants visible in these images explicitly agreed to be visualized.
Figure 5: Additional images of the data recording showing the variety of everyday activities our dataset includes.
14 Initial signal verification
In Fig. 6, we show the raw mean intensity values after spatial cropping of the skin region and the eye region (see Fig. 4) compared to the ground truth contact PPG signal from the nose. We can clearly see that the blood volume pulse is present both in the eyes and skin region with the skin region having a higher signal-to-noise ratio (SNR) compared to the eyes.
Figure 6: Example raw mean intensity of the skin and eye region, showing the higher SNR for the skin region around the eyes compared to the eyes.
15 Hyperparameter optimization
In Tab. 9, we show the results of EgoPulseFormer when training our model with different hyperparameters such as different image sizes, input window sizes, or target signals.
Hyperparameter Configuration MAE RMSE MAPE r
Image preprocessing Raw 16.33 20.04 20.93 0.31
Standardized 14.44 18.21 18.66 0.46
Difference 8.99 12.09 10.96 0.80
Standardized difference 8.82 12.03 10.82 0.81
Image size $24\times 64$ 9.73 12.69 12.12 0.77
$\boldsymbol{48\times 128}$ 8.82 12.03 10.82 0.81
$96\times 256$ 10.27 13.10 12.46 0.79
Window size 64 9.95 12.82 11.93 0.80
128 8.82 12.03 10.82 0.81
256 9.71 13.02 11.41 0.78
Loss function Negative Pearson [92] 10.62 13.57 13.16 0.81
MAE 9.71 13.02 11.41 0.78
MSE 8.82 12.03 10.82 0.81
Target signal Heart rate 10.99 14.03 13.49 0.77
Nose PPG 8.82 12.03 10.82 0.81
Augmentation Without 10.57 14.19 12.78 0.70
With 8.82 12.03 10.82 0.81
Table 9: Results of our model on the egoPPG-DB dataset when training with different hyperparameters.